# Walk Wisely on Graph: Knowledge Graph Reasoning with Dual Agents via Efficient Guidance-Exploration

### v1

@@ Introduction
Over the past few years, several embedding-based models (Bordes et al. 2013; Lin et al. 2015) are proposed to focus on preserving the structural information in the KG and are effective for single-hop reasoning, but lack the necessary interpretability. To solve this problem, some works (Xiong, Hoang, and Wang 2017; Das et al. 2018) have considered introducing reinforcement learning framework to KG reasoning tasks, these approaches compose single-hop triplets into multi-hop reasoning chains. With the development of deep learning (DL) and reinforcement learning (RL), recent multi-hop works (Wang et al. 2019; Lv et al. 2020; Nikopensius et al. 2023; Xu et al. 2024) incorporate more advanced DL and RL techniques. For example, AttnPath (Wang et al. 2019) makes use of attention mechanism to force an agent to walk forward every step to avoid the agent stalling at the same entity node constantly.
Although existing multi-hop reasoning models have achieved impressive results, there is a noteworthy issue with these models that they only perform well when the agent’s reasoning path length is short. The drawback that the agent relies heavily on short reasoning chains is fatal in certain datasets, such as sparse KG. Figure 1 is an illustration of short direct path and long indirect path. It’s challenging for an agent to identify a short direct path just like (a) due to the sparsity which makes the relation $r_{2}$ do not exist. The optimal path for agent is $e_{1}\rightarrow e_{2}\rightarrow e_{4}\rightarrow e_{3}$ we called long indirect path instead of $e_{1}\rightarrow e_{2}\rightarrow e_{3}$. From this point, enhancing the agent’s long-distance reasoning ability is one solution to alleviate the multi-hop models’ poor performance in the sparse KG.
Figure 1: An illustrative example of (a) short direct path and (b) long indirect path. When no short direct path exists, the agent searches for a long indirect path.
However, the dimension of the discrete action space at each space is large (Das et al. 2018). As the path length increases, the choices the agent faced will grow exponentially. The most relevant to our work are CURL (Zhang et al. 2022) which proposed a dual-agent framework and mutual reinforcement rewards to assign one of the agents (GIANT) searching on cluster-level paths quickly and providing stage-wise hints for another agent (DWARF). Compared to the classic multi-hop model MINERVA (Das et al. 2018), CURL does improve long-distance reasoning ability. But CURL also has two drawbacks, which make CURL’s performance inconsistent. (1) Mutual reinforcement reward mechanism forces the low-level agent DWARF to adopt similar policies to the high-level agent GIANT, even if GIANT’s policies is not well enough. This may result in false-negative rewards to the intermediate actions which are reasonable. (2) At the early phase, DWARF and GIANT adopt a near-random policy which makes low training efficacy. For briefness, we call the high-level agent as GIANT and the low-level agent as DWARF like (Bai et al. 2022).
A number of efforts have been made to justify Transformer in the RL architecture.
@@
In line with the focus of our work, we provide a brief overview of the background and related work on knowledge graph embedding and multi-hop reasoning.
@@
Knowledge graph embedding (KGE) methods map entities to vectors in low-dimensional embedding space, and model relations as transformations between entity embeddings (Bai et al. 2022). Once we map them to low-dimensional dense vector space, we can use modeling methods to perform calculations and reasoning (Ma et al. 2018). Prominent examples include TransE (Bordes et al. 2013), TransR (Lin et al. 2015), ConvE (Dettmers et al. 2018) and TuckER (Balazevic, Allen, and Hospedales 2019). These models are equipped with a scoring function that maps any triplet $(e_{s},r_{q},e_{o})$ to a scalar score. For instance, TransE (Bordes et al. 2013) first encoded the entities and relations into latent vectors by following translational principle in point-wise Euclidean space. The translation assumption in TransE implies the equation that $||\mathbf{h}+\mathbf{r}-\mathbf{t}||\approx 0$. The embeddings of entities and relations are learnt by maximizing the score function so that the likelihood score is high for true triplets while low for false triplets.
@@
The advancement of deep reinforcement learning (DRL) has led to an increased interest in utilizing DRL to address path-finding issues. The first groundbreaking work combining DRL and KG reasoning is DeepPath (Xiong, Hoang, and Wang 2017), which has an important inspiration for the subsequent derivative models. However, the method is constrained by the requirement to know the target entity in advance. MINERVA (Das et al. 2018), by contrast, does not need to know the target entity in progress but instead causes the agent to walk on the knowledge graph until it finds answer.
In recent years, many powerful neural networks have been developed. Some works consider the potential of utilizing these networks to generate walking policies. M-Walk (Shen et al. 2018) utilizes RNN to record the sequence trajectories obtained by the agent traversal and then optimize the discounted reward by a RL algorithm called Monte Carlo Tree Search (MCTS), to tackle the problem of sparse rewards. GRL (Wang et al. 2020) employs both generative adversarial net (GAN) and long short-term memory (LSTM) to generate new trajectory sequences if needed, which means the agent cannot only reason in the original graph, but sub-graphs can be automatically generated to extend the relational path until the target entities are found. SQUIRE (Bai et al. 2022) utilizes an encoder-decoder Transformer structure to map query and previous path sequence to a contextualized representation, and further use such representation to autoregressively decode the output path, token by token. These models have demonstrated satisfactory performance in short-distance reasoning scenarios. However, their performance in long-distance reasoning tasks has consistently been suboptimal.
The proposal of hierarchical reinforcement learning (HRL) (Li et al. 2018; Wen et al. 2020; Liu et al. 2023) has brought attention to the fact that mupltiple agents can walk on the graph simultaneously. CURL (Zhang et al. 2022) aligns with the RL formulation of MINERVA (Das et al. 2018), i.e., learning to walk and search for answer entities of a particular KG query in an end-to-end fashion. Nevertheless, previous multi-agent based KG reasoning models don’t strike a satisfying balance between guidance from other agents and self-exploration. Moreover, these models don’t utilize enough of entity embedding and relation embedding, which wastes some structural information of the KG. In contrast to the preceding methodologies, FULORA employs a distinctive training approach that seamlessly integrates supervised learning and reinforcement learning. This innovative approach effectively reconciles the inherent trade-off between self-exploration and path-guidance, a challenge commonly encountered by DWARF.
@@
As in the standard reinforcement learning paradigm, we divide the environment which is modeled as a Markov decision process (MDP) into four elements: state, action, transition and reward.
State. The state consists of two entities and query relation. Specifically, the two entities consist of a current entity $e_{t}$ and a source entity $e_{s}$. The current entity visited by the agent at time step $t$ can be considered as state-dependent information, it reflects the reasoning condition of the agent. The source entity and query relation are the global context shared by all states. Concretely, $s_{t}=(e_{t},e_{s},r_{q})\in\mathcal{S}$.
Action. In reinforcement learning, the agent interacts with the environment by taking action. Specific to the graph walking task, the agent makes an action $a_{t}\in\mathcal{A}$ by selecting an outgoing edge of graph $\mathcal{G}$. This process can be formed as $a_{t}=\{(r_{t+1},e_{t+1})|(e_{t},r_{t+1},e_{t+1})\in\mathcal{G}\}$. In order to grant the agent an option to terminate a search, the agent can choose to stay at each entity by adding a self-loop edge to every action.
Transition. Choosing a new action entails a transition of state. A transition function $\delta:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is defined by $\delta(s_{t},a_{t})=\delta((e_{t},e_{s},r_{q}),a_{t})$, which means that performing an action at a given state produces a new state. Most multi-hop reasoning picks a neighboring entity at random because it’s common for an entity to have multiple neighbors that are all connected to it via the same relation (Xiong, Hoang, and Wang 2017; Das et al. 2018).
Reward. In the default formulation, after the agent takes an action under a certain state if the corresponding entity is a correct target, a favorable reward $1$ is obtained, and conversely, an unfavorable reward 00 is obtained.
@@
Entity Embeddings and Cluster Embeddings
Consistent with (Zhang et al. 2022), we employ TransE (Bordes et al. 2013) due to its efficiency in encoding the structural proximity information of KG to generate pre-trained entity embeddings, then we divide the original KG into $N$ clusters by utilizing K-means. In order for GIANT can easily walk on the cluster-level graph $\mathcal{G}^{c}$ while preserving the relation information of the original KG $\mathcal{G}$, we add a link to two clusters if there is at least one entity-level edge between them, named cluster mapping. An example of cluster mapping is shown in Figure 2.
Figure 2: An example of cluster mapping.
Policy Networks For GIANT and DWARF
In our model, we utilize a three-layer LSTM, enabling the agent to memorize and learn from the actions taken before. In contrast to previous models, the necessity arises to design the network separately in this case, given that two agents are walking on the KG. An agent based on LSTM encodes the recursive sequence as a continuous vector $\mathbf{h}_{t}$. Specifically, the hidden state embedding of GIANT is $\mathbf{h}^{c}_{t}$ while the hidden state embedding of DWARF is $\mathbf{h}^{e}_{t}$. Their initial hidden state is $\mathbf{0}$. In addition, we define an information sharing vector $\mathbf{I}_{t}=[\mathbf{h}^{e}_{t};\mathbf{h}^{c}_{t}]$ for GIANT and DWARF to share path information. To a certain extent, cluster-level paths are complementary to entity-level paths, as they ensure the sharing of essential path information from GIANT to DWARF.
@@
where $\mathbf{W}^{c}\in\mathbb{R}^{2d\times 4d}$ is a projection matrix to maintain shape.
For DWARF. Different from GIANT walking on the cluster KG, DWARF encounters greater difficulties on the original KG. Firstly, an entity has several different aspects. For example, a university professor may be linked with professional relations like worksForUniversity or worksInState, and also family relations like spouse or father/mother. Secondly, It can be observed that cluster-level paths are typically shorter than entity-level paths. Consequently, DWARF must possess enhanced long-distance reasoning capabilities to identify an accurate target. Furthermore, for diverse query relations, it’s advisable for DWARF to prioritize relations and neighbors that are highly correlated with the query relations. So we introduce Graph Attention mechanism (Velickovic et al. 2017) into DWARF.
A single-layer feedforward neural network is employed to compute attention weight. Consider that the current entity is $e_{i}$, a collection of all neighbors is $\mathcal{N}(i)$. The attention weight from $e_{i}$ and its neighbor $e_{j}\in\mathcal{N}(i)$ is calculated as
$\alpha_{ij}=\text{LeakyReLu}(\mathbf{a}[\mathbf{W}\mathbf{e}_{i}^{\top};\mathbf{W}\mathbf{e}_{j}^{\top}]),$ (2)
where $\mathbf{W}\in\mathbb{R}^{d\times d}$ is a linear transformation matrix and $\mathbf{a}\in\mathbb{R}^{2d}$ is a weight vector shared across all entities. For $e_{i}$, we normalize the attention weights with SoftMax:
$\alpha_{ij}=\frac{e^{\alpha_{ij}}}{\sum_{k\in\mathcal{N}(i)}e^{\alpha_{ik}}}.$ (3)
Now we can compute the graph attention for the state:
$\mathbf{atn}_{i}=\sum_{k\in\mathcal{N}(i)}\alpha_{ik}\mathbf{W}\mathbf{e}_{k}.$ (4)
DWARF’s history embedding $\mathbf{h}_{t}^{e}$ can be obtained from
$\mathbf{h}_{t}^{e}=\text{LSTM}_{e}(\mathbf{W}^{e}[\mathbf{h}_{t-1}^{e};\mathbf{atn}_{t-1};\mathbf{I}_{t-1}]),\mathbf{a}_{t-1}^{e}),$ (5)
where $\mathbf{W}^{e}\in\mathbb{R}^{2d\times 5d}$ is a projection matrix, while the action representation $\mathbf{a}_{t}^{e}$ is the concatenation of the relation embedding $\mathbf{r}_{t}\in\mathbb{R}^{d}$ and the end node embedding $\mathbf{e}_{t}\in\mathbb{R}^{d}$, i.e., $\mathbf{a}_{t}^{e}=[\mathbf{r}_{t};\mathbf{e}_{t}]\in\mathbb{R}^{2d}$.
@@
$\begin{array}[]{ll}&\mathbf{d}^{c}_{t}=\text{SoftMax}(\mathbf{A}_{t}^{c}\times\mathbf{W}_{2}^{c}\text{ReLU}(\mathbf{W}^{c}_{1}[\mathbf{c}_{t};\mathbf{h}_{t}^{c}])),\\ &a_{t}^{c}\sim\text{Categorical}(\textbf{d}^{c}_{t}),\end{array}$ (6)
$\begin{array}[]{ll}&\mathbf{d}^{e}_{t}=\text{SoftMax}(\mathbf{A}_{t}^{e}\times\mathbf{W}_{2}^{e}\text{ReLU}(\mathbf{W}^{e}_{1}[\mathbf{e}_{t};\mathbf{r}_{q};\mathbf{h}_{t}^{e}])),\\ &a_{t}^{e}\sim\text{Categorical}(\textbf{d}^{e}_{t}),\end{array}$ (7)
where $\mathbf{W}_{1}^{c},\mathbf{W}_{2}^{c}\in\mathbb{R}^{4d\times 4d}$ and $\mathbf{W}_{1}^{e},\mathbf{W}_{2}^{e}\in\mathbb{R}^{6d\times 6d}$ are the matrices of learnable weights to maintain dimension of history embedding. While $\mathbf{A}_{t}^{c}\in\mathbb{R}^{|A_{t}^{c}|\times 4d},\mathbf{A}_{t}^{e}\in\mathbb{R}^{|A_{t}^{e}|\times 6d}$ represent the embeddings of all next possible actions for GIANT and DWARF respectively.
@@
As mentioned above, in order to alleviate the issue of sparse rewards and huge action space, we hope GIANT will guide DWARF to shrink the action space. Nevertheless, the guidance provided by GIANT is not always beneficial, primarily due to two key issues: (1) Poor guidance will force DWARF to get wrong answer, and (2) even if GIANT is capable of generating fair guidance, mapping from entity to cluster results in a distribution shift, that is, a appropriate policy for GIANT may not be fully appropriate for DWARF.
@@
Constraint Reward.  Considering the constraint, that is, receiving high-quality guidance from GIANT, we introduce a metric for the state distance between GIANT and DWARF, denoted as $D(s_{t}^{c},s_{t}^{e})$. It is calculated as the cosine similarity between the pre-trained embeddings of the current entity and the current cluster:
$D(s_{t}^{c},s_{t}^{e})=\frac{\mathbf{c}_{t}^{\top}\mathbf{e}_{t}}{||\mathbf{c}_{t}||_{2}||\mathbf{e}_{t}||_{2}}.$ (8)
@@
maximize $\text{maximize }\mathbb{E}_{a_{1}^{e},...,a_{T}^{e}\sim\pi_{\theta}^{e}}\left[\sum_{t=0}^{T-1}r_{e}(s_{t}^{e}|s_{0}^{e})\right]$ (9)
subject to $\text{subject to }D(s_{t}^{c},s_{t}^{e})\geq\frac{\delta}{r_{c}(s_{t}^{c})+\varepsilon}$
Figure 4: An overview of FULORA framework.
where $r_{e}(s_{t}^{e})$ and $r_{c}(s_{t}^{c})$ are default rewards for DWARF and GIANT respectively, the agent obtains a favorable reward 1 if the corresponding entity or cluster is a correct target and unfavorable reward 0 otherwise. It is only when GIANT reaches the correct cluster that the constraint in Equation 9 is operative.
@@
$J(\theta_{\pi^{e}})=\sum_{t=0}^{T-1}[(1-\lambda(s_{t}^{e}))r_{e}(s_{t}^{e})+\lambda(s_{t}^{e})D(s_{t}^{c},s_{t}^{e})],$ (10)
the Lagrange multiplier is updated to make the guidance-exploration trade-off. When the state similarity metric between GIANT and DWARF $D(s_{t}^{c},s_{t}^{e})$ is less than the threshold, the parameter $\theta_{\lambda}$ are optimized so that $\lambda(s_{t}^{e})$ increases to encourage GIANT to provide more guidance for DWARF. Conversely, the parameter $\theta_{\lambda}$ are optimized so that $\lambda(s_{t}^{e})$ decreases to encourage DWARF to explore in the constrictive space when the state similarity metric between GIANT and DWARF $D(s_{t}^{c},s_{t}^{e})$ is greater than the threshold. In practice, we utilize a cross-entropy loss to optimize $\theta_{\lambda}$:
@@
$\text{where }y=\left\{\begin{aligned} &0\quad\quad\text{if}\ D(s_{t}^{c},s_{t}^{e})\geq\frac{\delta}{r_{c}(s_{t}^{c})+\varepsilon}\\ &1\quad\quad\text{otherwise}\end{aligned}\right.$
@@
A further challenge is to enhance the search efficiency of GIANT in order to provide DWARF with the requisite guidance as expeditiously as possible, given that the search is limited to a fixed number of step $T$. In the default reward, GIANT will receive 1 reward only when it reaches the correct cluster, and the rest will be 0, which makes GIANT adopt random policy at the early phase. This phenomenon is not conducive to stable learning outcomes for two agents. On the basis of the theory of reward shaping (Ng, Harada, and Russell 1999; Harutyunyan et al. 2015), we rewrite the reward function of GIANT, named dynamic path feedback. In particular, we write it in the form of an objective function $J(\theta_{\pi^{c}})$,

### v3

@@ Introduction
Over recent years, embedding-based models (Bordes et al. 2013; Lin et al. 2015) have effectively preserved KG structural information for single-hop reasoning but lack interpretability. To address this, reinforcement learning (RL) frameworks (Xiong, Hoang, and Wang 2017; Das et al. 2018) have been introduced to compose single-hop triplets into multi-hop reasoning chains. Recent advances in deep learning (DL) and RL (Wang et al. 2019; Lv et al. 2020; Nikopensius et al. 2023; Wang et al. 2025) further enhance multi-hop reasoning. For instance, AttnPath (Wang et al. 2019) employs attention mechanisms to guide agents, preventing them from stalling at the same node.
Figure 1: An illustrative example of short direct path and long indirect path. When no short direct path exists, the agent searches for a long indirect path.
Although existing multi-hop reasoning models have achieved impressive results, there is a noteworthy issue with these models that they only perform well when the agent’s reasoning path length is short. The drawback that the agent relies heavily on short reasoning chains is fatal in certain datasets, such as sparse KG. Figure 1 is an illustration of short direct path and long indirect path. It’s challenging for an agent to identify a short direct path due to the sparsity which makes the relation $r_{2}$ do not exist. The optimal path for agent is $e_{1}\rightarrow e_{2}\rightarrow e_{4}\rightarrow e_{3}$ we called long indirect path instead of $e_{1}\rightarrow e_{2}\rightarrow e_{3}$. From this point, enhancing the agent’s long-distance reasoning ability is one solution to alleviate the multi-hop models’ poor performance in the sparse KG.
However, the dimension of the discrete action space at each space is large (Das et al. 2018). As the path length increases, the choices the agent faced will grow exponentially. The most relevant to our work is CURL (Zhang et al. 2022) which proposed a dual-agent framework and mutual reinforcement rewards to assign one of the agents (GIANT) searching on cluster-level paths quickly and providing stage-wise hints for another agent (DWARF). Compared to the classic multi-hop model MINERVA (Das et al. 2018), CURL does improve long-distance reasoning ability. But CURL also has two drawbacks, which make CURL’s performance inconsistent: (1) Mutual reinforcement reward mechanism forces the low-level agent DWARF to adopt similar policies to the high-level agent GIANT, even if GIANT’s policies is not well enough. This may result in false-negative rewards to the intermediate actions which are reasonable. (2) At the early phase, DWARF and GIANT adopt a near-random policy which makes low training efficacy. For briefness, we call the high-level agent as GIANT and the low-level agent as DWARF like (Bai et al. 2022).
@@
In line with the focus of our work, we provide a brief overview of the background and related work on knowledge graph reasoning.
@@
Knowledge graph embedding (KGE) methods map entities to vectors in low-dimensional embedding space, and model relations as transformations between entity embeddings (Bai et al. 2022). Once we map them to low-dimensional dense vector space, we can use modeling methods to perform calculations and reasoning (Ma et al. 2018). Prominent examples include TransE (Bordes et al. 2013), TransR (Lin et al. 2015), ConvE (Dettmers et al. 2018), TuckER (Balazevic, Allen, and Hospedales 2019) and LMKE (Wang et al. 2022), which are equipped with a scoring function that maps any triplet $(e_{s},r_{q},e_{o})$ to a scalar score. While KGE is effective at capturing simpler relationships in the graph, such as first-order adjacency relationships, it struggles with more complex reasoning tasks, particularly those involving multi-hop reasoning (Yao, Mao, and Luo 2019; Wang et al. 2023).
GNN-based Reasoning
Graph neural networks (GNNs) (Scarselli et al. 2008; Veličković et al. 2017; Xu et al. 2018) are a class of models used for representation learning, specifically designed to encode the structural information of graphs. In the context of link prediction, common frameworks often rely on an auto-encoder formulation, where GNNs generate node embeddings, and edges are predicted as a function of node pairs. These frameworks are inductive when node features are provided, but they become transductive when such features are not available. Another set of frameworks, including SEAL (Zhang and Chen 2018) and GraIL (Teru, Denis, and Hamilton 2020), explicitly encode the subgraph surrounding each node pair for link prediction. Recent advancements in this field include works such as NBFNet (Zhu et al. 2021) and RED-GNN (Zhang and Yao 2022). The former solves the path formulation with learned operators in the generalized Bellman-Ford algorithm while the latter makes use of dynamic programming to recursively encodes multiple r-digraphs with shared edges, and utilizes query-dependent attention mechanism to select the strongly correlated edges. However, these methods primarily focus on the structure of the knowledge graph (KG) itself, specifically the graph structure information of the central node, without addressing how to enhance the ability for long-distance reasoning.
@@
The advancement of deep reinforcement learning (DRL) has sparked interest in applying DRL to path-finding tasks. The first significant work combining DRL and KG reasoning is DeepPath (Xiong, Hoang, and Wang 2017), which inspired subsequent models, though it requires prior knowledge of the target entity. In contrast, MINERVA (Das et al. 2018) eliminates this requirement, allowing the agent to traverse the knowledge graph until it finds the target. Recent developments have explored the use of powerful neural networks for generating walking policies. M-Walk (Shen et al. 2018) uses an RNN (Elman 1990) to record agent trajectories and optimize rewards with a Monte Carlo Tree Search (MCTS) (Coulom 2006). GRL (Wang et al. 2020) combines GAN (Goodfellow et al. 2014) and LSTM (Hochreiter 1997) to generate new trajectory sequences, enabling the agent to reason not only within the original graph but also in automatically generated sub-graphs, extending relational paths until target entities are found. HMLS (Zheng et al. 2024) improves the generalizability and effectiveness of multi-hop reasoning in few-shot scenarios by exploiting hard relations and hierarchical relation structures. While these models perform well in short-distance reasoning (path length = 3), their performance in long-distance reasoning tasks remains suboptimal. Unlike previous multi-hop reasoning methods, FULORA focuses more on the agents’ long-distance reasoning ability and the significant impact of sparse rewards on training efficiency.
Figure 2: An overview of FULORA framework. ❶ Given a KG $\mathcal{G}$, We first pre-embed the KG using TransE and then apply K-means clustering to generate the cluster-level KG $\mathcal{G}^{c}$. ❷ We design separate policy networks for GIANT and DWARF, using cluster-level KG $\mathcal{G}^{c}$ and entity-level KG $\mathcal{G}^{e}$ as inputs. The hidden states of GIANT and DWARF $\mathbf{h}_{t}^{c}$ and $\mathbf{h}_{t}^{e}$ share information, facilitating communication. ❸ To enable DWARF to better leverage the KG structure, we apply the graph attention mechanism and feed the resulting attention vector into the policy network. Dynamic Path Feedback alleviates the near-random policy issue caused by sparse rewards in early training phase, allowing GIANT to provide high-quality guidance to DWARF sooner.
@@
State. The state consists of two entities—current entity $e_{t}$ and source entity $e_{s}$—and the query relation. The current entity $e_{t}$ is state-dependent, reflecting the agent’s reasoning condition, while the source entity $e_{s}$ and query relation are shared globally. Formally, $s_{t}=(e_{t},e_{s},r_{q})\in\mathcal{S}$.
Action. The agent interacts with the environment by selecting an action $a_{t}\in\mathcal{A}$, corresponding to an outgoing edge of the graph $\mathcal{G}$. Formally, $a_{t}=\{(r_{t+1},e_{t+1})|(e_{t},r_{t+1},e_{t+1})\in\mathcal{G}\}$. To allow termination, the agent can also stay at the current entity by selecting a self-loop edge.
Transition. A state transition occurs when the agent takes an action. The transition function $\delta:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ is defined as $\delta(s_{t},a_{t})$, which produces a new state. Multi-hop reasoning typically selects a neighboring entity randomly, as entities often have multiple neighbors connected by the same relation (Xiong, Hoang, and Wang 2017; Das et al. 2018).
Reward. After the agent takes an action, if the corresponding entity is the correct target, it receives a reward of 1; otherwise, it receives a reward of 0.
@@
Embedding Generation
Consistent with (Zhang et al. 2022), we employ TransE (Bordes et al. 2013) due to its efficiency in encoding the structural proximity information of KG to generate pre-trained entity embeddings, then we divide the original KG into $N$ clusters by utilizing K-means. In order for GIANT can easily walk on the cluster-level graph $\mathcal{G}^{c}$ while preserving the relation information of the original KG $\mathcal{G}$, we add a link to two clusters if there is at least one entity-level edge between them, which is detailed in Appendix E [^1] Our extended version is available at
https://arxiv.org/pdf/2408.01880
Policy Networks
In our model, we utilize a three-layer LSTM, enabling the agent to memorize and learn from the actions taken before. In contrast to previous models, the necessity arises to design the network separately in this case, given that two agents are walking on the KG. An agent based on LSTM encodes the recursive sequence as a continuous vector $\mathbf{h}_{t}\in\mathbb{R}^{2d}$. Specifically, the hidden state embedding of GIANT is $\mathbf{h}^{c}_{t}$ while the hidden state embedding of DWARF is $\mathbf{h}^{e}_{t}$. Their initial hidden state is $\mathbf{0}$. In addition, we define an information sharing vector $\mathbf{I}_{t}=[\mathbf{h}^{e}_{t};\mathbf{h}^{c}_{t}]$ for GIANT and DWARF to share path information. To a certain extent, cluster-level paths are complementary to entity-level paths, as they ensure the sharing of essential path information from GIANT to DWARF.
@@
where $\mathbf{W}^{c}\in\mathbb{R}^{2d\times 6d}$ is a projection matrix to maintain shape.
For DWARF. Unlike GIANT, which operates on cluster-level KGs, DWARF faces greater challenges on the original KG. Entities often have multiple aspects. For instance, a professor may have both professional relations (e.g., worksForUniversity) and family relations (e.g., spouse). Additionally, cluster-level paths are typically shorter than entity-level paths, requiring DWARF to have enhanced long-distance reasoning. Consequently, DWARF should prioritize relations and neighbors most relevant to the query. Therefore, we integrate the Graph Attention mechanism (Velickovic et al. 2017) into DWARF. We adopt the same approach as AttnPath (Wang et al. 2019) to obtain the attention vector $\mathbf{atn}_{t-1}$. DWARF’s history embedding $\mathbf{h}_{t}^{e}$ can be obtained from
$\mathbf{h}_{t}^{e}=\text{LSTM}_{e}(\mathbf{W}^{e}[\mathbf{h}_{t-1}^{e};\mathbf{atn}_{t-1};\mathbf{I}_{t-1}],\mathbf{a}_{t-1}^{e}),$ (2)
where $\mathbf{W}^{e}\in\mathbb{R}^{2d\times 7d}$ is a projection matrix, while the action representation $\mathbf{a}_{t}^{e}$ is the concatenation of the relation embedding $\mathbf{r}_{t}\in\mathbb{R}^{d}$ and the end node embedding $\mathbf{e}_{t}\in\mathbb{R}^{d}$, i.e., $\mathbf{a}_{t}^{e}=[\mathbf{r}_{t};\mathbf{e}_{t}]\in\mathbb{R}^{2d}$.
@@
$\mathbf{d}^{c}_{t}=\text{SoftMax}(\mathbf{A}_{t}^{c}\times\mathbf{W}_{2}^{c}\text{ReLU}(\mathbf{W}^{c}_{1}[\mathbf{c}_{t};\mathbf{h}_{t}^{c}])),$ (3)
$a_{t}^{c}\sim\text{Categorical}(\textbf{d}^{c}_{t}),$
$\mathbf{d}^{e}_{t}=\text{SoftMax}(\mathbf{A}_{t}^{e}\times\mathbf{W}_{2}^{e}\text{ReLU}(\mathbf{W}^{e}_{1}[\mathbf{e}_{t};\mathbf{r}_{q};\mathbf{h}_{t}^{e}])),$ (4)
$a_{t}^{e}\sim\text{Categorical}(\textbf{d}^{e}_{t}),$
where $\mathbf{W}_{1}^{c}\in\mathbb{R}^{4d\times 4d}$, $\mathbf{W}_{2}^{c}\in\mathbb{R}^{2d\times 4d}$, $\mathbf{W}_{1}^{e}\in\mathbb{R}^{6d\times 6d}$, $\mathbf{W}_{2}^{e}\in\mathbb{R}^{2d\times 6d}$ are the matrices of learnable weights to maintain dimension of history embedding. While $\mathbf{A}_{t}^{c}\in\mathbb{R}^{|A_{t}^{c}|\times 2d},\mathbf{A}_{t}^{e}\in\mathbb{R}^{|A_{t}^{e}|\times 2d}$ represent the embeddings of all next possible actions for GIANT and DWARF respectively.
@@
As mentioned above, to address the issue of sparse rewards and a large action space, we aim for GIANT to guide DWARF in reducing the action space. However, GIANT’s guidance is not always beneficial due to two key issues:
  • Poor guidance. Poor guidance can lead DWARF to incorrect answers.
  • Policy shift. A policy which is suitable for GIANT may not be fully applicable to DWARF.
@@
Constraint Reward.  Considering the constraint, that is, receiving high-quality guidance from GIANT, we introduce a metric for the state similarity between GIANT and DWARF, denoted as $\text{Sim}(s_{t}^{c},s_{t}^{e})$. It is calculated as the cosine similarity between the pre-trained embeddings of the current cluster and the current entity:
$\text{Sim}(s_{t}^{c},s_{t}^{e})=\frac{\mathbf{c}_{t}^{\top}\mathbf{e}_{t}}{||\mathbf{c}_{t}||_{2}||\mathbf{e}_{t}||_{2}}.$ (5)
@@
max $\text{max }\mathbb{E}_{a_{1}^{e},...,a_{T}^{e}\sim\pi_{\theta}^{e}}\left[\sum_{t=0}^{T-1}r_{e}(s_{t}^{e}|s_{0}^{e})\right]$ (6)
[…]
where $r_{e}(s_{t}^{e})$ and $r_{c}(s_{t}^{c})$ are default rewards for DWARF and GIANT respectively, the agent obtains a favorable reward 1 if the corresponding entity or cluster is a correct target and unfavorable reward 0 otherwise. It is only when GIANT reaches the correct cluster that the constraint in Equation 6 is operative. In practice, we set $\varepsilon=-0.01\delta$.
@@
$J(\theta_{\pi^{e}})=\sum_{t=0}^{T-1}[(1-\lambda(s_{t}^{e}))r_{e}(s_{t}^{e})+\lambda(s_{t}^{e})\text{Sim}(s_{t}^{c},s_{t}^{e})],$ (7)
the Lagrange multiplier is updated to make the guidance-exploration trade-off. Figure 3 gives an illustration: When the state similarity metric between GIANT and DWARF $D(s_{t}^{c},s_{t}^{e})$ is less than the threshold, the parameter $\theta_{\lambda}$ are optimized so that $\lambda(s_{t}^{e})$ increases to encourage GIANT to provide more guidance for DWARF. Conversely, the parameter $\theta_{\lambda}$ are optimized so that $\lambda(s_{t}^{e})$ decreases to encourage DWARF to explore in the constrictive space when the state similarity metric between GIANT and DWARF $D(s_{t}^{c},s_{t}^{e})$ is greater than the threshold. In practice, we utilize a cross-entropy loss to optimize $\theta_{\lambda}$:
@@
$\text{where }y=\left\{\begin{aligned} &0\quad\quad\text{if}\ \text{Sim}(s_{t}^{c},s_{t}^{e})\geq\frac{\delta}{r_{c}(s_{t}^{c})+\varepsilon}\\ &1\quad\quad\text{otherwise}\end{aligned}\right.$
@@
A further challenge is to enhance the search efficiency of GIANT in order to provide DWARF with the requisite guidance as expeditiously as possible, given that the search is limited to a fixed number of step $T$. In the default reward, GIANT will receive 1 reward only when it reaches the correct cluster, and the rest will be 0, which makes GIANT adopt random policy at the early phase. This phenomenon is not conducive to stable learning outcomes for two agents. On the basis of the theory of reward shaping (Harutyunyan et al. 2015), we rewrite the reward function of GIANT, named dynamic path feedback. In particular, we write it in the form of an objective function $J(\theta_{\pi^{c}})$,